# SSD: A State-based Stealthy Backdoor Attack For Navigation System in UAV Route Planning

### v1

@@ 3.2 Detector
However, the system may generate outliers due to attacks, noise, faults, and other factors, all of which contribute to the measurements deviating from the true values. To prevent the EKF state from these disruptive outliers, implementing an outlier detection mechanism becomes crucial. The chi-square statistical test serves as an efficient tool for determining outliers [25, 26]. It evaluates the current measured value by calculating the chi-square statistic $\chi^{2}_{k}$, comparing it to a predefined statistical significance threshold. When $\chi^{2}_{k}$ surpasses this threshold $\tau$, the measurement is considered as an outlier, and suitable measures are undertaken, including discarding the measurement or executing a partial update. The chi-square statistic $\chi^{2}_{k}$ is defined as:
@@
4 Security Analysis
4.1 Attack Formulation
@@
4.2 Study of Attack
The GNSS observation matrix $H_{GNSS}$ are
@@ 4.2 Study of Attack
From Eq.16, we can see that the state transfer error and the measurement error affect the magnitude of the gain $K_{k}$ simultaneously. $Q_{k}$ and $R_{k}$ reflect the ability to cover systematic uncertainty and measurement uncertainty, respectively. Therefore, inappropriate $Q_{k}$ and $R_{k}$ can lead to filter divergence or biased estimation. However, in INS, $Q_{k}$ and $R_{k}$ are generally determined based on a priori knowledge by pre-running the filter calculations offline and remain constant during the filtering process online. As a result, both the process estimation error covariance $R_{k}$ and the Kalman gain $K_{k}$ converge quickly and remain constant during the process demonstrates that the value of $K_{k}$ is determined by the ratio of $Q_{k}$ and $R_{k}$.
@@
    Firstly, When performing maneuvers, the system is highly nonlinear. i.e.
@@
    In addition, this complex nonlinear characteristic will be further expanded due to physical factors such as the lag of IMU data and the presence of friction in the gyroscope. According to Eq. 8 and 9, the accumulation of linearisation errors will increase $e_{i}$ and thus increase the process noise $P_{i}$. leading to inaccurate mathematical modeling and huge nonlinear errors. The fixed $Q_{k}$ makes it difficult to suppress the nonlinear error increased due to the change of motion state. According to Eq. 16, the value of $K_{i}$ will indirectly increase, making the INS more inclined to trust the GNSS data. Therefore, in this scenario, the prediction of the system model cannot effectively reflect the actual physical process.
  • R uncertainty: The modification of GNSS results in a shift in the measurement data distribution, rendering a fixed R inadequate for accurately describing the measurement noise distribution. Consequently, delta experiences a significant increase. Additionally, the rising k value leads the system to place greater trust in GNSS, further amplifying the impact of the attack associated with delta.
We choose biased signal attack and multiplicative Attack to evaluate the phenomenon above. We let the UAV maintain a linear velocity of $5m/s$ and perform two typical motion modes, uniform linear motion (linear motion) and uniform circular motion (non-linear motion), respectively. For each flight state, we choose $2s$ for the GNSS attack respectively and observe the changes in the localization Error $LocErr$ in $6s$ before and after the attack. The experimental results are shown in Figure 3(a) and 3(b). Thus we can get Finding 1.
@@
Figure 3: Study of Attack
Finding 1: For a GNSS spoof of the same magnitude, applying it during the UAV’s non-linear motion results in more pronounced fluctuations in positioning accuracy compared to linear motion. These changes in motion dynamics create greater vulnerabilities for INS.
Position and velocity are tightly coupled in INS, making velocity attacks easier to detect. However, there has been limited research on attacks targeting velocity. We would like to explore one question: Are there obvious correlations between the velocity attack stealthiness and the UAV motion states? We select a 10-second flight trajectory under both linear and nonlinear motion states for velocity and positional attacks. We adopt a bias signal attack to evaluate stealthiness and observe the changes in the chi-square detector. Fig. 3(c) and 3(d) indicate that the detector shows no significant fluctuations between the two motion states for the positional attack. However, velocity attacks demonstrate a greater sensitivity to motion states, with significantly lower cardinality detector results observed in nonlinear motion states. This reveal key insights, summarized as Finding 2.
@@
5 Attack Design
From the analysis of Section 4.2, we observe that UAV exhibits greater vulnerability in a non-linear motion state compared to a linear one due to the combined effects of model and measurement uncertainty. During non-linear motion, an attack of the same magnitude can produce a more significant change in the navigation output than in the linear motion state. However, this vulnerability arises only when the uncertainty is heightened due to changes in the motion state. Moreover, Finding2 indicates that positional attack in non-linear motion will increase the risk of being detected. This introduces a key challenge to the attacker:
C1: How to opportunistically exploit these vulnerable periods to acheive maximum localization error while maintaining stealth.
To address C1, we design a backdoor-like attack to exploit these vulnerable periods directly. Inspired by the backdoor attacks in deep networks (DNNs) [6, 27, 28], we proposed a novel stated-based stealthy backdoor (SSD) attack against INS in a route planning scenario. SSD utilizes the motion state as a trigger, enabling the UAV triggers an attack in a nonlinear motion state while maintaining normal operation in a linear motion state. Based on this design, not only can the duration of the attack be greatly reduced, but also the mission completion rate of the UAV can be effectively reduced. (UAV mission completion is usually accompanied by large maneuvers.)
@@ 5 Attack Design
where $t_{i}$ is the attack time for attackers. $\theta$, $\alpha$ and $\phi$ are hyperparameters that can be dynamically adjusted to maintain an equilibrium between stealthiness and effectiveness. We will demonstrate how to configure these parameters in Section 6.2. SSDs only attack position when acceleration is 0 while velocity attacks are triggered when acceleration is non-zero. Therefore, SSD uses acceleration as a trigger. It uses organic coupling between velocity and position to perform a combined attack. Its pseudocode are presented in Algorithm 1.
Input: victim UAV position $P_{i}=(P_{Ni},P_{Ei},P_{Di})$, victim UAV velocity $V_{i}=(V_{Ni},V_{Ei},V_{Di})$, IMU sample frequency $F_{imu}$,
@@ 6      Receive victim UAV velocity $V_{i}$ and position $P_{i}$ from GNSS ;
7       calculate acceleration $a_{i}=(\frac{\partial V_{Ni}}{\partial t_{i}},\frac{\partial V_{Ei}}{\partial t_{i}},\frac{\partial V_{Di}}{\partial t_{i}})$;
@@ 15      Fuse $P_{i}$ and $V_{i}$ data to update $X_{i+1}$
6 Experiment
6.1 Experimental Setup
6.1.1 Trajectory Dataset
@@
  • Trajectory I: Straight-line Path As illustrated in Fig. 5(a), this trajectory is designed for executing simple tasks, such as flying to a specific location and performing actions along a predefined path (e.g. patrols, cargo transport, etc.).
@@
6.1.2 Navigation Algorithm
We choose the estimation and control library EKF (ECL EKF2) of PX4 drone autopilot [29] project as the target navigation algorithms
  • ECL EKF2 implements EKF to estimate pose in the NED reference frame by fusing MARG (magnetic, angular rate, gravity) and GPS data. MARG data is derived from magnetometer, gyroscope, and accelerometer sensors. It uses a 22-element state vector to track the orientation quaternion, velocity, position, MARG sensor biases, and geomagnetic.
@@ 6.1.2 Navigation Algorithm
Accelerometers and gyroscopes operate at relatively high sample rates and necessitate high-rate processing. In contrast, GNSS and magnetometers function at relatively low sampling rates and require lower data processing rates. To replicate this configuration in our experiment, the IMUs (accelerometers, gyroscopes, and magnetometers) were sampled at 160 Hz, while the GNSS was sampled at 1 Hz. As demonstrated in Fig 6, only one out of every 160 samples from the magnetometer was provided to the fusion algorithm.
6.1.4 Evaluation Metircs
We utilize three metrics to evaluate the effectiveness of SSD comprehensively. Initially, we include two metrics that are widely used in relevant studies.
@@ 6.1.4 Evaluation Metircs
However, the two indicators above alone are insufficient to fully capture the impact of targeted attacks on UAVs. This introduces another challenge for evaluating SSD:
@@
where $P_{i}^{a}$ and $P_{i}$ refers to the predicted positions before and after the attack, respectively. $N_{w}$ refers to the number of waypoints. APDE provides a more nuanced understanding of how targeted attacks impact the UAV’s ability to adhere to its prescribed flight path, particularly at critical waypoints, thus enabling a more accurate assessment of the potential risks to mission success.
6.2 Parametric Analysis
@@
The effectiveness and stealthiness of the attack are highly dependent on the choice of parameters. We analyze the sensitivity of SSD to various parameters by combining different values for $\theta$, $\alpha$, and $\phi$ across multiple scenarios. To measure the attack effectiveness and stealthiness, we use ADE and the maximum chi-square statistic (denoted as $\chi^{2}_{max}$), respectively. The experimental results are shown in Fig 7. In positional attacks, as $\theta$ increases, $\chi^{2}_{max}$ exhibits a corresponding upward trend. We further observe that the rate of increase in $\chi^{2}_{max}$ slows as $\alpha$ increases. This suggests that higher values of $\alpha$ help mitigate the growth of $\chi^{2}_{max}$. Notably, when $\alpha$ exceeds a critical threshold value 13, $\chi^{2}_{max}$ starts to stabilize, converging within a relatively narrow threshold range. Within this range, $\chi^{2}_{max}$ fluctuates minimally and remains largely unaffected by changes in $\theta$, demonstrating high stability and consistency. For the velocity attack, as shown in Fig. 7(b), $\chi^{2}_{max}$ gradually converges to about 6 when $\phi$ is less than 0.08.
6.3 Ablation study
To validate the effectiveness of velocity-based and position-based attacks, we conducted an experiment where the UAV performed a 35-s flight incorporating linear and nonlinear motion states. The first 20 seconds involved uniform linear motion at a velocity of $2m/s$. Afterward, the UAV transitioned into a uniform circular motion mode, maintaining a linear velocity of $2m/s$ for the remaining 15 seconds. This flight trajectory was used for the ablation experiments.
6.3.1 Contributions of different attacks
To explore the contribution of velocity and positional attacks to overall attack effectiveness, we designed comparison experiments with three attack strategies: single position perturbation attack (SPA), single velocity perturbation attack (SVA), and combined concerted attack (CCA). The quantitative analysis of ADEs, presented in Table II, shows that both individual attacks are effective. Specifically, the ADEs for SPA and SVA are improved by 184% and 212%, respectively, compared to the baseline. When comparing the combined attack (CCA) to the single attacks, the ADE improves by 129% over SPA and 112% over SVA. This indicates that CCA not only significantly enhances attack effectiveness by leveraging the coupling effect between position and velocity, but also perturbs both longitude and latitude directions simultaneously. Furthermore, the stealth assessment results in Fig. 8 show that while the CCA approach induces brief oscillations in the detection statistics, these fluctuations remain well within the acceptable threshold limits, ensuring that the attack remains stealthy.
TABLE II: Ablation Study of Different Attack’s Contribution
Position N Position E
@@ 6.3.1 Contributions of different attacks
Figure 8: $\chi^{2}$ comparation
6.3.2 Attack Combination
We apply the velocity attack during the linear motion state and the position attack during the nonlinear motion state and examine the impact of this combination on the attack’s stealthiness. The experimental results, shown in Fig. 9, reveal that $9s$ after the attack begins, the detector quickly identifies the existence of the attack. Furthermore, even when the UAV transitions to a nonlinear motion state, the detector continues to successfully detect the attack for $5s$, despite the change in attack mode. This phenomenon occurs because the drastic velocity changes in the linear motion state cause significant fluctuations in the residuals. In contrast, the nonlinear motion state allows SSD to effectively smooth out the impact of the velocity attack, ensuring the attack remains stealthy throughout the process. This further validates the rationale of the SSD framework.
@@ 6.3.2 Attack Combination
6.4 Attack Stealthiness
To validate the stealthiness, we selected chi-square detector for attack detection. We also chose biased signal and multiplicative attacks for comparison and evaluated them across three mission trajectories. Based on prior research, we designed the following attack payloads: (1) GNSS positions are added with a uniform distribution U(0, 0.0005); (2) GNSS positions are scaled by a factor of 1.5. The experimental results in Fig. 10 demonstrate that SSD successfully limits detection statics to the threshold range and bypasses both detection methods with a carefully chosen set of attack parameters. Notably, compared to these works, we significantly reduced the loadings applied for the biased signal Attack. However, both detection methods were able to detect the attack effectively. It is clear that SSD exhibits strong stealthiness properties under chi-square detection methods and several motion states.

### v2

@@ III-B Detector
However, the system may generate outliers due to attacks, noise, faults, and other factors, all of which contribute to the measurements deviating from the true values. To prevent the EKF state from these disruptive outliers, implementing an outlier detection mechanism becomes crucial. The chi-square statistical test serves as an efficient tool for determining outliers [35, 36]. It evaluates the current measured value by calculating the chi-square statistic $\chi^{2}_{k}$, comparing it to a predefined statistical significance threshold. When $\chi^{2}_{k}$ surpasses this threshold $\tau$, the measurement is considered an outlier, and suitable measures are undertaken, including discarding the measurement or executing a partial update. The chi-square statistic $\chi^{2}_{k}$ is defined as:
@@
IV Security Analysis
IV-A Attack Formulation
@@
IV-B Study of Attack
The GNSS observation matrix $H_{GNSS}$ are as follows.
@@ IV-B Study of Attack
From Eq.16, we can see that the state transfer error and the measurement error affect the magnitude of the gain $K_{k}$ simultaneously. $Q_{k}$ and $R_{k}$ reflect the ability to cover systematic uncertainty and measurement uncertainty, respectively. Therefore, inappropriate $Q_{k}$ and $R_{k}$ can lead to filter divergence or biased estimation. However, in INS, $Q_{k}$ and $R_{k}$ are generally determined based on a priori knowledge by pre-running the filter calculations offline and remain constant during the filtering process online. As a result, both the process estimation error covariance $R_{k}$ and the Kalman gain $K_{k}$ converge quickly and remain constant during the process, demonstrating that the value of $K_{k}$ is determined by the ratio of $Q_{k}$ and $R_{k}$.
@@
    Firstly, when performing maneuvers, the system is highly nonlinear. i.e.,
@@
    According to Eq. 8 and 9, the accumulation of linearisation errors will increase $e_{i}$ and thus increase the process noise $P_{i}$. leading to inaccurate mathematical modeling and huge nonlinear errors. The fixed $Q_{k}$ makes it difficult to suppress the nonlinear error increased due to the change of motion state. According to Eq. 16, the value of $K_{i}$ will indirectly increase, making the INS more inclined to trust the GNSS data. In addition, this complex nonlinear characteristic will be further expanded due to physical factors such as the lag of IMU data and the presence of friction in the gyroscope. Therefore, in this scenario, the prediction of the system model cannot effectively reflect the actual physical process.
  • R uncertainty: The modification of GNSS results in a shift in the measurement data distribution, rendering a fixed $R_{k}$ inadequate for accurately describing the measurement noise distribution. Consequently, $\delta_{i}$ experiences a significant increase. Additionally, the rising $K_{k}$ value leads the system to place greater trust in GNSS, further amplifying the impact of the attack associated with $\delta_{i}$.
We employ a biased signal attack and a multiplicative attack to evaluate the phenomenon above. The UAV maintains a constant velocity of $5m/s$ and performs two typical motion modes, uniform linear motion (linear motion) and uniform circular motion (non-linear motion), respectively. For each flight state, we apply an attack window of two attack inputs for the GNSS respectively and observe the changes in the localization Error $LocErr$ before and after the attack. As a result, the attack time is 2 seconds since the GNSS input is 1 Hz. The experimental results are shown in Fig. 3(a) and 3(b). Thus, we can get Finding 1.
@@
Figure 3: Study of GNSS Attack under Different Motion states
Finding 1: For a GNSS spoofing of the same magnitude, applying it during the UAV’s non-linear motion results in more pronounced fluctuations in positioning accuracy compared to linear motion. These changes in motion dynamics create greater vulnerabilities for INS.
Position and velocity are tightly coupled in INS, making velocity attacks easier to modify the position result. However, there has been limited research on attacks targeting velocity. We would like to explore one question: Are there obvious correlations between the velocity attack stealthiness and the UAV motion states? To validate this question, we select two flight trajectories with linear and nonlinear motion states respectively. We use a biased signal attack to perturb velocity and position measurement and evaluate stealthiness by observing the changes in the chi-square detector. Fig. 3(c) and 3(d) indicate that the detector shows no significant fluctuations between the two motion states for the positional attack. However, velocity attacks demonstrate a greater sensitivity to motion states, with significantly lower cardinality detector results observed in nonlinear motion states. This reveals key insights, summarized as Finding 2.
@@
V Attack Design
From the analysis of Section IV-B, we observe that UAV exhibits greater vulnerability in a non-linear motion state compared to a linear one due to the combined effects of model and measurement uncertainty. During non-linear motion, an attack of the same magnitude can produce a more significant change in the navigation output than in the linear motion state. However, this vulnerability arises only when the uncertainty is heightened due to changes in the motion state. Moreover, Finding2 indicates that a positional attack in non-linear motion will increase the risk of being detected. This introduces a key challenge to the attacker:
C1: How to opportunistically exploit these vulnerable periods to achieve maximum localization error while maintaining stealth.
To address C1, we design a backdoor-like attack to exploit these vulnerable periods directly. Inspired by the backdoor attacks in deep networks (DNNs) [12, 37, 38], we proposed a novel stated-based stealthy backdoor (SSD) attack against INS in a route planning scenario. SSD utilizes the motion state as a trigger, enabling the UAV to trigger an attack in a nonlinear motion state while maintaining normal operation in a linear motion state. Based on this design, not only can the detection rates of the attack be greatly reduced, but also the mission completion rate of the UAV can be effectively reduced. (UAV mission completion is usually accompanied by large maneuvers.)
@@ V Attack Design
where $t_{i}$ is the attack time for attackers. $\theta$, $\alpha$, and $\phi$ are hyperparameters that can be dynamically adjusted to maintain an equilibrium between stealthiness and effectiveness. We will demonstrate how to configure these parameters in Section VI-B. SSD chooses acceleration as a trigger. It uses organic coupling between velocity and position to perform a combined attack. Its pseudocode is presented in Algorithm 1.
Input: Victim UAV position $P_{i}=(P_{Ni},P_{Ei},P_{Di})$, Victim UAV velocity $V_{i}=(V_{Ni},V_{Ei},V_{Di})$, IMU sampling frequency $F_{imu}$, Iteration number $M$
@@ 6      Receive victim UAV velocity $V_{i}$ and position $P_{i}$ from GNSS ;
7       Compute acceleration $a_{i}=(\frac{\partial V_{Ni}}{\partial t_{i}},\frac{\partial V_{Ei}}{\partial t_{i}},\frac{\partial V_{Di}}{\partial t_{i}})$;
@@
9             // Apply position perturbation in linear motion state ;
@@ 10             $P_{i}=P_{i}+\theta{e^{ti/\alpha}}$;
13             // Apply velocity perturbation in nonlinear motion state ;
@@ 17      Fuse $P_{i}$ and $V_{i}$ to update $X_{i+1}$
VI Experiment
VI-A Experimental Setup
VI-A1 Trajectory Dataset
@@
  • Trajectory I: Straight-line Path As illustrated in Fig. 5(a), this trajectory is designed for executing simple tasks, such as flying to a specific location and performing actions along a predefined path (e.g., patrols, cargo transport, etc.).
@@
VI-A2 Navigation Algorithm
We choose the estimation and control library EKF (ECL EKF2) of the PX4 drone autopilot [39] project as the target navigation algorithms.
  • ECL EKF2 implements EKF to estimate pose in the NED reference frame by fusing MARG (magnetic, angular rate, gravity) and GNSS data. MARG data is derived from magnetometer, gyroscope, and accelerometer sensors. It uses a 22-element state vector to track the orientation quaternion, velocity, position, MARG sensor biases, and geomagnetic.
@@ VI-A2 Navigation Algorithm
Accelerometers and gyroscopes operate at relatively high sample rates and necessitate high-rate processing. In contrast, GNSS and magnetometers function at relatively low sampling rates and require lower data processing rates. To replicate this configuration in our experiment, the IMUs (accelerometers, gyroscopes, and magnetometers) were sampled at 160 Hz, while the GNSS was sampled at 1 Hz. As demonstrated in Fig 6, only one out of every 160 samples from the IMUs was provided to the fusion algorithm.
VI-A4 Evaluation Metircs
We utilize three metrics to evaluate the effectiveness of SSD comprehensively. Initially, we include two metrics that are widely used in relevant studies [40].
@@ VI-A4 Evaluation Metircs
However, the two indicators above alone are insufficient to capture the impact of targeted attacks on UAVs. This introduces another challenge for evaluating SSD:
@@
where $P_{i}^{a}$ and $P_{i}$ refer to the predicted positions before and after the attack, respectively. $N_{w}$ refers to the number of waypoints. APDE provides a more nuanced understanding of how targeted attacks impact the UAV’s ability to adhere to its prescribed flight path, particularly at critical waypoints, thus enabling a more accurate assessment of the potential risks to mission success.
VI-B Parametric Analysis
@@
The effectiveness and stealthiness of the attack are highly dependent on the choice of parameters. We analyze the sensitivity of SSD to various parameters by combining different values for $\theta$, $\alpha$, and $\phi$ across multiple scenarios (the same as Section IV-B). To measure the attack effectiveness and stealthiness, we use ADE and the maximum chi-square statistic (denoted as $\chi^{2}_{max}$), respectively. The experimental results are shown in Fig 7. In positional attacks, as $\theta$ increases, $\chi^{2}_{max}$ exhibits a corresponding upward trend. We further observe that the rate of increase in $\chi^{2}_{max}$ slows as $\alpha$ increases. This suggests that higher values of $\alpha$ help mitigate the growth of $\chi^{2}_{max}$. Notably, when $\alpha$ exceeds a critical threshold value 13, $\chi^{2}_{max}$ starts to stabilize, converging within a relatively narrow threshold range. Within this range, $\chi^{2}_{max}$ fluctuates minimally and remains largely unaffected by changes in $\theta$, demonstrating high stability and consistency. For the velocity attack, as shown in Fig. 7(b), $\chi^{2}_{max}$ gradually converges to about 6 when $\phi$ is less than 0.08. In the following experiments, we set the values of $\theta$, $\alpha$, and $\phi$ to 20, 11, and 0.08, respectively.
VI-C Ablation study
To validate the effectiveness of velocity-based and position-based attacks, we conducted an experiment where the UAV performed a 35 second flight incorporating linear and nonlinear motion states. The first 20 seconds involved uniform linear motion at a velocity of $2m/s$. Afterward, the UAV transitioned into a uniform circular motion mode, maintaining a linear velocity of $2m/s$ for the remaining 15 seconds. This flight trajectory was used for the ablation experiments.
VI-C1 Contributions of different attacks
To explore the contribution of velocity and positional attacks to overall attack effectiveness, we designed comparison experiments with three attack strategies: single position perturbation attack (SPA), single velocity perturbation attack (SVA), and combined concerted attack (CCA). The quantitative analysis of ADEs, presented in Table II, shows that both individual attacks are effective. Specifically, the ADEs for SPA and SVA are improved by 184% and 212%, respectively, compared to the baseline. When comparing the combined attack (CCA) to the single attacks, the ADE improves by 129% over SPA and 112% over SVA. This indicates that CCA not only significantly enhances attack effectiveness by leveraging the coupling effect between position and velocity but also perturbs both longitude and latitude directions simultaneously. Furthermore, the stealth assessment results in Fig. 8 show that while the CCA approach induces brief oscillations in the detection statistics, these fluctuations remain well within the acceptable threshold limits, ensuring that the attack remains stealthy.
TABLE II: Ablation Study of Different Attacks’ Contribution
Position N (meters) Position E (meters)
@@ VI-C1 Contributions of different attacks
Figure 8: Stealthiness Comparison for Attack Contribution
VI-C2 Attack Combination
We apply the velocity attack during the linear motion state and the position attack during the nonlinear motion state and examine the impact of this combination on the attack’s stealthiness. The experimental results, shown in Fig. 9, reveal that $9s$ after the attack begins, the detector quickly identifies the existence of the attack. Furthermore, even when the UAV transitions to a nonlinear motion state, the detector continues to successfully detect the attack for $5s$, despite the change in attack mode. This phenomenon occurs because the drastic velocity changes in the linear motion state cause significant fluctuations in the residuals. In contrast, the nonlinear motion state allows SSD to effectively smooth out the impact of the velocity attack, ensuring the attack remains stealthy throughout the process. This result, along with Finding2, further validates the rationale of the SSD framework.
@@ VI-C2 Attack Combination
VI-D Attack Stealthiness
To validate the stealthiness, we selected a chi-square detector for attack detection. For the threshold, we selected a value corresponding to a 95% confidence level, which is 11.1. We also chose biased signal and multiplicative attacks for comparison and evaluated them across three mission trajectories. Based on prior research, we designed the following attack payloads: (1) GNSS positions are added with a uniform distribution U(0, 0.0005); (2) GNSS positions are scaled by a factor of 1.5. The experimental results in Fig. 10 demonstrate that SSD successfully limits detection statistics to the threshold range and bypasses both detection methods with a carefully chosen set of attack parameters. Notably, compared to these works [5, 41], we significantly reduced the loadings applied for the biased signal Attack. However, both detection methods were able to detect the attack effectively. It is clear that SSD exhibits strong stealthiness properties under chi-square detection methods and several motion states.